OpenStudy (anonymous):

Part 1: Is the following equation true? (2 points) Part 2: Use complete sentences to explain the properties used in making your decision. (6 points) 2 log 5 x + 1/2 log 5 y – 4 log 5 z = log 5 x^2√y/z^4.

OpenStudy (campbell_st):

there are all base 5 logs the equation is \[2\log _{5}(x)+1/2\log _{5}(y) - 4\log _{5}(z)\]

OpenStudy (anonymous):

ok. Then its true

OpenStudy (campbell_st):

the equation is true log laws are identical to power laws... log law for powers of a term \[\log _{a}(b^y) = ylog _{a}(b)\] this has been applied to all three terms next.. log laws for products add the terms in the product \[\log _{a}(bc) = \log _{a}(b) +\log _{a}(c)\] log laws for division.... subtract the terms \[\log _{a}(\frac{b}{c}) = \log _{a}(b) - \log _{a}(c)\]

OpenStudy (anonymous):

thanks dude :)